# Convergence of patterned matrices with random walk entries

### v1

@@ 2.4 Inverse stable subordinator and fractional Poisson process
For more details on different characterizations of the fPP, we refer the reader to [19, 21, 22], and references therein.
@@ 3 Patterned random matrices with CTRW entries
Motivated by all the discussions so far, let $\{X_{i,j,k},i\geq 1,\,j\geq 1,\,k\geq 1\}$ be a triple index sequence of real valued rvs. Let us consider a collection of CTRWs $\{\{S_{i,j,n}(t)\}_{t\geq 0}\}_{1\leq i,j\leq n}$, $n\in\mathbb{N}$ defined as follows:
$S_{i,j}^{(n)}(t)\coloneqq\sum_{k=1}^{[nt]}X_{i,j,k},\ 1\leq i,j\leq n,\ n\in\mathbb{N}.$ (3.1)
Now we focus on certain large dimensional patterned random matrices whose entries are from (3.1), via an associated link function $L$, taking values in either $\mathbb{N}$ or $\mathbb{N}^{2}$. When $L$ takes values in $\mathbb{N}$, by a slight abuse of notation, we regard the triple indexed sequence $\{X_{i,j,k},\ i\geq 1,\,j\geq 1,\,k\geq 1\}$ as a bi-indexed sequence $\{X_{i,k},\ i\geq 1,\,k\geq 1\}$. Let us consider the sequence $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ of matrix valued random processes with link function $L$ defined as follows:
$A_{n}(t)=(S_{L(i,j)}^{(n)}(t))_{1\leq i,j\leq n},\ n\geq 1,\ t\geq 0.$ (3.2)
$A_{n}(0)$ is taken to be the $n\times n$ matrix of zeroes. Consider the scaled matrix $n^{-1}A_{n}(t)$, with eigenvalues are $\lambda_{1}(t),\dots,\lambda_{n}(t)$. Its ESD is given by $F_{n^{-1}A_{n}(t)}(x)\coloneqq n^{-1}\sum_{r=1}^{n}\mathbb{I}(\lambda_{r}(t)\leq x),\ x\in\mathbb{R}$, $t>0$. From (2.2), the $l$th moment of its EESD is
$\frac{\mathbb{E}\mathrm{Trace}(A_{n}(t))^{l}}{n^{1+l}}=\sum_{w:\ \text{$w$ is word of length $l$}}\frac{1}{n^{1+l}}\sum_{w\in\Pi(w)}\mathbb{E}S_{\pi}^{(n)}(t),\ l\geq 1,\ t\geq 0,$ (3.3)
where $S_{\pi}^{(n)}(t)=S_{L(\pi(0),\pi(1))}^{(n)}(t)\dots S_{L(\pi(l-1),\pi(l))}^{(n)}(t)$.
@@
Assumption I Let $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ be a triple indexed sequence of independent rvs with zero mean and unit variance such that $\sup_{i,j,k}\mathbb{E}|X_{i,j,k}|^{l}<K_{l}<\infty$ for all $l\geq 3$.
3.1 Marginal convergence
Using Assumption I, the random variables $n^{-1/2}\sum_{k=1}^{n}X_{ijk}$ are independent and have uniformly bounded moments of all orders, a crucial assumption in Theorem 7.5.1 of [9]. The following two theorems follow from that theorem, and hence we omit their proofs.
@@ 3.1 Marginal convergence
Let $\{X_{i,j,k},\ 1\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ be a triple indexed sequence that satisfies Assumption I. Let $A_{n}(t)$ be the random matrix as defined in (3.2) with symmetric link function $L$ that satisfies Property B. Assume that the following limits exist:
@@
where $\Pi(w)$ denotes the cardinality of the equivalence class $\Pi(w)$. Then, the limiting values of moments defined in (3.3) are given by
@@
$\mathbb{E}\big(n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}-\mathbb{E}n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}\big)^{4}=t^{2l}O(n^{-2}),\ l\geq 1,\ t\geq 0.$
Thus, (M4) holds and the ESD of $n^{-1}A_{n}(t)$ converges weakly a.s. surely to the unique probability law with moments as in (3.5).
@@
Let $A_{n}^{(1)},A_{n}^{(2)},\dots,A_{n}^{(n)}$ be independent $n\times n$ random matrices with input sequences $\{X_{i,j,1}\},\{X_{i,j,2}\},\dots,\{X_{i,j,n}\}$, respectively, each satisfying Assumption I, and a common symmetric link function $L$, satisfying Property B. If the limits (3.4) exist, then for each $j=1,\dots,n$, the EESD of $B_{n}^{(j)}=n^{-1/2}A_{n}^{(j)}$ weakly converges to a probability law whose $l$th moment equals $\mu^{(l)}(1)$ for every $l\geq 1$, defined in (3.5). Moreover, this convergence holds true for their ESD a.s. (see [9]).
From Theorem 3.1 and Theorem 3.2 for $t=1$, it follows that the ESD of $n^{-1/2}(B_{n}^{(1)}+\dots+B_{n}^{(n)})$ also converges weakly a.s. to the probability law whose $l$th moment is given by $\mu^{(l)}(1)$ for all $l\geq 1$. Thus, $B_{n}^{(1)},\dots,B_{n}^{(n)}$ and $n^{-1}(A_{n}^{(1)}+\dots+A_{n}^{(n)})$ have same LSD.
@@
Let $\{X_{i,j,k},\ 1\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ be a triple indexed sequence that satisfies Assumption I. Let $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ be a sequence of matrix valued random processes as defined in (3.2) with link function $L$. Then,
(i) for $L$ defined by $L(i,j)=(\max\{i,j\},\min\{i,j\})$, the EESD of $n^{-1}A_{n}(t)$ converges to the semi-circular law with zero mean and variance $t$;
(ii) If L(i,j)=|ij|L(i,j)=|i-j| then the EESD of $n^{-1}A_{n}(t)$ converges to a probability law whose corresponding $(2m)$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iii) If L(i,j)=i+j2L(i,j)=i+j-2 then the EESD of $n^{-1}A_{n}(t)$ converges to a probability law, symmetric about zero, whose corresponding $2m$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iv) If L(i,j)=i+j2modnL(i,j)=i+j-2\ $\text${mod}\ n, $1\leq i,j\leq n$ then the EESD of $n^{-1}A_{n}(t)$ converges to a symmetric Rayleigh law whose $(2m)$th moment is given by $t^{m}m!$, $m\geq 1$;
(v) If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$ then the EESD of $n^{-1}A_{n}(t)$ converges to the Gaussian law with mean zero and variance $t$.
3.2 Joint convergence

### v2

@@ 2.2 Inverse stable subordinator and fractional Poisson process
For details on different characterizations of the fPP, see [19, 21, 22], and references therein.
2.3 Free Brownian motion (FBM) and time-change
(i) $B(0)=0$;
A Brownian motion time-changed by an independent inverse stable subordinator is a key mathematical model for analyzing anomalous diffusion in statistical mechanics. It is used for the modeling of motion of a particle experiencing a trapping effect.
Here, we define a random time-changed FBM driven by an inverse stable subordinator. Let $\{B(t)\}_{t\geq 0}$ be an FBM on the $W^{*}$-probability space $(\mathcal{W},\tau)$. Let $L^{\alpha}(t),t\geq 0$ be an inverse stable subordinator on $(\Omega,\mathcal{F},\mathbb{P})$, independent of the FBM. From [31], all moments of $L^{\alpha}(t),t\geq 0$ are finite, Hence $\mathcal{A}_{\infty}$, the set of all polynomials in $L^{\alpha}(t),t\geq 0$ is an NCP with its state as expectation operator $\mathbb{E}$. Let us consider a collection $\{B^{\alpha}(t)\}_{t\geq 0}$ of variables of the tensor product space $(\hat{\mathcal{A}}:=\mathcal{W}\otimes\mathcal{A}_{\infty},\hat{\tau}:=\mathbb{E}\tau)$ defined as follows:
$B^{\alpha}(t)\coloneqq$ $B(L^{\alpha}(t)),\ 0<\alpha\leq 1,$ (2.3)
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l}))\coloneqq$ $\mathbb{E}\tau(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l}))$
for all $l\geq 1$ and $t_{1},\dots,t_{l}\geq 0$. Equivalently,
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l}))=\int_{\mathbb{R}^{l}_{+}}\tau(B(x_{1})\dots B(x_{l}))\mathrm{Pr}\{L^{\alpha}(t_{1})\in\mathrm{d}x_{1},\dots,L^{\alpha}(t_{l})\in\mathrm{d}x_{l}\}.$
$\{B^{\alpha}(t)\}_{t\geq 0}$ will be called a time-changed FBM. For $\alpha=1$, it reduces to the FBM.
where $C_{m}=(2m)!/(m+1)!m!$ is the $m$th Catalan number.
Proposition 2.1.
Let $Z(t)$ be a semi-circle rv with mean zero and variance $t>0$ that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$. Then,
$\mathbb{E}e^{uZ(L^{\alpha}(t))}=(u^{2}t^{\alpha})^{-1}\Big(W_{\alpha,1-\alpha}(u^{2}t^{\alpha})-\frac{1}{\Gamma(1-\alpha)}\Big),\ 0<\alpha<1,\ u\in\mathbb{R},\ t>0,$ (2.5)
where $W_{\nu,\mu}$, $\nu,\mu>0$ is the Wright function (for definition see [18], Eq. (1.11.1)). As a consequence, $\{B^{\alpha}(t)\}_{t\geq 0}$ is a free process with $B^{\alpha}(0)=0$ and for each $t>0$, the spectral distribution of $B^{\alpha}(t)$ has a valid probability density with support $\mathbb{R}$, whose associated moment generating function (mgf) is given by (2.5).
Proof.
2.4 Patterned random matrices and link function
Note that the notion of convergence as defined in Section 2.1.2 applies to $n\times n$ random matrices with the choice of state being $n^{-1}\mathbb{E}\text{Trace}$. A related concept of convergence is that of the spectral distribution. Let $A_{n}$ be an $n\times n$ symmetric random matrix whose entries are rvs on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and its eigenvalues are $\lambda_{1},\dots,\lambda_{n}$. The empirical spectral distribution (ESD) of $A_{n}$ is defined as follows:
A class of patterned random matrices that includes many important matrices has been studied in a unified way by [10]. We introduce their basic ideas now. For $d\in\{1,2\}$, let us consider a sequence $\{L^{d}_{n}\}_{n\geq 1}$ of functions defined as $L^{d}_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ with $L^{d}_{n+1}(i,j)=L^{d}_{n}(i,j)$. By using a common notation $L^{d}\coloneqq L^{d}_{n}$, it is called the link function.
Let $\{a_{i},\ i\geq 1\}$ and $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ be sequence and bi-sequence of rvs. For $n\geq 1$, a $n\times n$ random matrix $A_{n}$ with link function $L^{d}$ and input sequence $\{a_{i},\ i\geq 1\}$ or $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ is defined as $A_{n}\coloneqq((a_{L^{d}(i,j)}))_{1\leq i,j\leq n}$. The following are examples of link functions for some well known random matrices:
(i) $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$: Wigner matrix;
(ii) $L^{1}(i,j)=|i-j|$: (symmetric) Toepliz matrix;
(iii) $L^{1}(i,j)=i+j-2$: (symmetric) Hankel matrix;
(iv) $L^{1}(i,j)=i+j-2\ \text{mod}\ n$, $1\leq i,j\leq n$: Reverse Circulant matrix;
(v) $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$: Symmetric Circulant matrix.
The following property that all the above link functions satisfy, plays a crucial role:
Property B
$\Delta L^{d}\coloneqq\sup_{n}\sup_{r\in\mathbb{Z}^{d}_{+}}\sup_{1\leq i\leq n}\#\{j:1\leq j\leq n,\ L^{d}(i,j)=r\}<\infty$.
From (2.6), the $l$th moment associated to the ESD of $A_{n}$ is $\mu^{(l)}_{F_{A_{n}}}=n^{-1}\mathrm{Trace}(A_{n}^{l})$, $l\geq 1$, and the $l$th moment associated to its EESD is given by the trace-moment formula: $\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=n^{-1}\mathbb{E}\mathrm{Trace}(A_{n}^{l})$. Let us now consider a non random sequence $\{\mu^{(l)}\}_{l\geq 1}$ such that for all $l\geq 1$, $\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\rightarrow\mu^{(l)}$ as $n\rightarrow\infty$. If the sequence $\{\mu^{(l)}\}_{l\geq 1}$ satisfies Carleman’s condition, $\sum_{l=1}^{\infty}(\mu^{(2l)})^{-1/2l}=\infty$, then $\{\mu^{(l)}\}_{l\geq 1}$ uniquely determines a probability distribution, and in that case the ESD converges in probability to this distribution. The convergence can be strengthened to a.s. sense if $\sum_{l=1}^{\infty}\mathbb{E}\big(\mu^{(l)}_{F_{A_{n}}}-\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\big)^{4}<\infty$ for all $n\geq 1$.This route has been taken for patterned random matrices. For example, see [9, 10, 11] and references therein, for several convergence results for patterned random matrices, both for the ESD and for joint convergence.
2.4.1 Matched circuits and words
For $l\geq 1$ and $n\geq 1$, a function $\pi:\{0,1,\dots,l\}\rightarrow\{1,2,\dots,n\}$ with $\pi(0)=\pi(l)$ is called a circuit of length $l\geq 1$. A circuit $\pi$ is called matched if all the $L^{d}$-values $L^{d}(\pi(j-1),\pi(j))$ are repeated. It is called pair-matched if each value $L^{d}(\pi(j-1),\pi(j))$ appears exactly twice. We can group these circuits into equivalence classes: circuits $\pi_{1}$ and $\pi_{2}$ belong to the same class, denoted as $\pi_{1}\sim\pi_{2}$ if $L^{d}(\pi_{1}(j-1),\pi_{1}(j))=L^{d}(\pi_{1}(i-1),\pi_{1}(i))$ if and only if $L^{d}(\pi_{2}(j-1),\pi_{2}(j))=L^{d}(\pi_{2}(i-1),\pi_{2}(i))$ for all $1\leq i,j\leq l$. These equivalence classes can be indexed by partitions of $\{1,\dots,l\}$, which are represented by words $w$ of length $l$ in which letters appear in alphabetical order. Let $w(i)$ denote the $i$th letter of $w$ and $\Pi(w)$ be the equivalence class associated with word $w$. Then,
$\Pi(w)\coloneqq\{\pi:w(i)=w(j)\ \text{if and only if}\ L^{d}(\pi(j-1),\pi(j))=L^{d}(\pi(i-1),\pi(i))\}.$ (2.7)
A word $w$ is called matched if each letter repeats, and it is called pair-matched if each letter repeats exactly twice. Note that a pair-matched word is always of even length. We denote the set of pair-matched words of length $2m$, $m\geq 1$ by $\mathcal{P}_{2}(2m)$.
2.4.2 Joint convergence of independent random matrices
We now outline the idea of joint convergence of independent symmetric patterned matrices using the trace-moment formula. Let $A_{r,n}$, $r=1,2,\dots,p$ be $n\times n$ independent random matrices with common link function $L^{d}$. The $(i,j)$th entry of $A_{r,n}$ will be denoted by $A_{r,n}(L(i,j))$. Note that a typical monomial $g$ in terms of $\{A_{r,n}\}_{1\leq r\leq p}$ has the following form: $g(\{A_{r,n}\}_{1\leq r\leq p})=A_{r_{1},n}\dots A_{r_{l},n}$, where $r_{j}\in\{1,2,\dots,p\}$ for each $j=1,2,\dots,l$. The collection $\{A_{r,n}\}_{1\leq r\leq p}$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges jointly if for every fixed monomial $g$, the following converges:
$\tau_{n}(g)=n^{-1}\sum_{\pi:\,\text{$\pi$ is a circuit of length $m$}}\mathbb{E}A_{r_{1},n}(L^{d}(\pi(0),\pi(1)))\dots A_{r_{l}n}(L^{d}(\pi(l-1),\pi(l))).$
It turns out that monomials $g$ with odd number of factors always yield the limit 0 and hence without loss we assume that the order of $g$ is $2m$ for some integer $m$. A circuit $\pi$ of length $l$ is called index-matched if the value $(r_{j},L^{d}(\pi(j-1),\pi(j)))$ repeats for each $j=1,2,\dots,l$. Note that circuits are index-matched with reference to the monomial $g$. Let $I=\{\pi:\ \pi\ \text{is a index-matched circuit}\}$. Consider a equivalence class on $I$ where for any two circuits $\pi_{1}$ and $\pi_{2}$, we have $\pi_{1}\sim\pi_{2}$ if for any $1\leq i,j\leq l$ such that $r_{i}=r_{j}$, $A_{r_{i},n}(L^{d}(\pi_{1}(i-1),\pi_{1}(i)))=A_{r_{j},n}(L^{d}(\pi_{1}(j-1),\pi_{1}(j)))$ if and only if $A_{r_{i},n}(L^{d}(\pi_{2}(i-1),\pi_{2}(i)))=A_{r_{j},n}(L^{d}(\pi_{2}(j-1),\pi_{2}(j)))$. Any such equivalence class can be indexed by a partition of $\{1,2,\dots,l\}$, which further induces an indexed word $w$ of length $l$, where the indexed letters appear in alphabetical order. An indexed word is called pair-matched if every indexed letter in it appears exactly twice. Only such letters remain relevant in the limit. Let $w_{r_{j}}(j)$ denote the $j$th letter of indexed matched word $w$ and $\Pi_{g}(w)$ denote the class of circuits corresponding to $w$ and monomial $g$. Then,
$\Pi_{g}(w)=\{\pi:\ w_{r_{i}}(i)=w_{r_{j}}(j)\ \text{iff}\ A_{r_{i},n}(L^{d}(\pi(i-1),\pi(i)))=A_{r_{j},n}(L^{d}(\pi(j-1),\pi(j)))\}.$
Using the results for a single matrix, this would show the joint convergence of the iid copies of the pattern matrices described above. For further details, see [6, 9, 10], and references therein.
@@ 3 Patterned random matrices with CTRW entries
Motivated by all the discussions so far, let $\{X_{i,j,k},i\geq 1,\,j\geq 1,\,k\geq 1\}$ and $\{X_{i,k},i\geq 0,\,k\geq 1\}$ be sequences of real valued rvs. Let us consider the collection of CTRWs $\{\{S_{i,j,n}(t)\}_{t\geq 0}\}_{1\leq i,j\leq n}$ and $\{\{S_{i,n}(t)\}_{t\geq 0}\}_{1\leq i,j\leq n}$, $n\in\mathbb{N}$ defined as follows:
$S_{i,j}^{(n)}(t)\coloneqq$ $\sum_{k=1}^{[nt]}X_{i,j,k},\ 1\leq i,j\leq n,\ n\in\mathbb{N},$ (3.1)
$S_{i}^{(n)}(t)\coloneqq$ $\sum_{k=1}^{[nt]}X_{i,k},\ 0\leq i\leq n,\ n\in\mathbb{N}.$ (3.2)
Let us consider the sequence $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ of matrix valued random processes with link function $L^{d}$ defined as follows:
$A_{n}(t)=((S_{L^{d}(i,j)}^{(n)}(t)))_{1\leq i,j\leq n},\ n\geq 1,\ t\geq 0.$ (3.3)
$A_{n}(0)$ is taken to be the $n\times n$ matrix of zeroes. Depending on whether $d=1$ or $d=2$, the entries of $A_{n}(t)$ are from the sequences in (3.1) or (3.2) respectively. Consider the scaled matrix $n^{-1}A_{n}(t)$, with eigenvalues are $\lambda_{1}(t),\dots,\lambda_{n}(t)$. Its ESD is given by $F_{n^{-1}A_{n}(t)}(x)\coloneqq n^{-1}\sum_{r=1}^{n}\mathbb{I}(\lambda_{r}(t)\leq x),\ x\in\mathbb{R}$, $t>0$.
@@
Assumption I $\{X_{i,j,k},X_{i,k},\ i\geq 0,\ j\geq 0,\ k\geq 1\}$ are independent rvs with zero mean and unit variance such that $\sup_{i,j,k}\mathbb{E}[|X_{i,j,k}|^{l}+|X_{i,k}|^{l}]<K_{l}<\infty$ for all $l\geq 3$.
3.1 Marginal algebraic convergence
Under Assumption I, the random variables $n^{-1/2}\sum_{k=1}^{n}X_{i,j,k}$ and $n^{-1/2}\sum_{k=1}^{n}X_{i,k}$ are independent with mean 0 and variance $1$. Moreover, the following bound holds for $S_{i,j}^{(n)}(t)$ and $S_{i}^{(n)}(t)$ defined in (3.1) and (3.2).
$\sup_{i,j,n,\ 0\leq t\leq K}\mathbb{E}[|n^{-1/2}S_{i,j}^{(n)}(t)|^{k}+|n^{-1/2}S_{i}^{(n)}(t)|^{k}]\leq K_{k}<\infty\ \ \text{for all integers}\ \ k\geq 1.$ (3.4)
This bound will be repeatedly without further mention. Due to (3.4), using Theorem 7.5.1 of [9], we immediately get the following:
@@ 3.1 Marginal algebraic convergence
Suppose $\{X_{i,j,k},\ X_{i,k},\ 0\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ satisfies Assumption I. Let $A_{n}(t)$ be the random matrix as defined in (3.3) with symmetric link function $L^{d}$ that satisfies Property B. Assume that the following limits exist:
@@
where $\Pi(w)$ denotes the cardinality of the equivalence class $\Pi(w)$ as defined in (2.7). Then,
@@
$\mathbb{E}\big(n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}-\mathbb{E}n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}\big)^{4}=t^{2l}O(n^{-2}),\ l\geq 1.$
Thus, the ESD of $n^{-1}A_{n}(t)$ converges weakly a.s. to a unique probability law with moments as in (3.6).
@@
Fix a symmetric link function $L^{d}$ that satisfying Property B. Let $A_{1,n},A_{2,n},\dots$ be independent $n\times n$ random matrices with link function $L^{d}$ and variables $\{X_{i,j,1}\},\{X_{i,j,2}\},\dots$ or $\{X_{i,1}\},\{X_{i,2}\},\dots$, depending on whether $d=1$ or $d=2$. Suppose the variables satisfy Assumption I. If the limits (3.5) exist, then for each $j\geq 1$, the EESD of $B_{j,n}=n^{-1/2}A_{j,n}$ weakly converges to a probability law whose $l$th moment equals $\mu^{(l)}(1)$ for every $l\geq 1$, defined in (3.6). Moreover, this convergence holds true for their ESD a.s. (see [9]).
From Theorem 3.1 and Theorem 3.2 for $t=1$, it follows that the ESD of $n^{-1/2}(B_{1,n}+\dots+B_{n,n})$ also converges weakly a.s. to the probability law whose $l$th moment is given by $\mu^{(l)}(1)$ for all $l\geq 1$. Thus, each $B_{j,n}$ and $n^{-1}(A_{1,n}+\dots+A_{n,n})$ have same LSD.
@@
Suppose $\{X_{i,j,k},\ X_{i,k},\ 1\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ satisfies Assumption I. Let $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ be as defined in (3.3) with link function $L^{d}$. Then, for each $t$,
(i) for $L^{2}$ defined by $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$, the ESD of $n^{-1}A_{n}(t)$ converges a.s. to the semi-circular law with zero mean and variance $t$;
(ii) If $L^{1}(i,j)=|i-j|$ then the ESD of $n^{-1}A_{n}(t)$ converges a.s. to a probability law whose corresponding $(2m)$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iii) If $L^{1}(i,j)=i+j-2$ then the ESD of $n^{-1}A_{n}(t)$ converges a.s. to a probability law, symmetric about zero, whose corresponding $2m$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iv) If $L^{1}(i,j)=i+j-2\ \text{mod}\ n$, $1\leq i,j\leq n$ then the ESD of $n^{-1}A_{n}(t)$ converges a.s. to a symmetric Rayleigh law whose $(2m)$th moment is given by $t^{m}m!$, $m\geq 1$;
(v) If $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$ then the ESD of $n^{-1}A_{n}(t)$ converges a.s. to the Gaussian law with mean zero and variance $t$.
3.2 Joint algebraic convergence